# Modeling and computation of the effective elastic behavior of parallelogram origami metamaterials

### v2

@@ 3 Modeling framework
Building on the theory, we now develop a constitutive model well-adapted to simulate the effective behavior of soft modes in parallelogram origami. A basic challenge is that it is hard to numerically implement exact geometrical constraints, like the ones in $A_{\text{eff}}$. This is especially true under prototypical displacement and traction boundary conditions on the effective deformation $\mathbf{y}$. We seek instead a practical continuum constitutive model that is faithful to the theory but also efficient to simulate and predictive.
@@ 3.1 Constitutive model
The division by $\sqrt{\det\mathbf{I}(\mathbf{y})}$ in the $W_{1}(\theta,\mathbf{I}(\mathbf{y}))$ seeks to prevent interpenetration of matter in the model, while the division by $|\tilde{\mathbf{u}}_{0}|^{4}|\tilde{\mathbf{v}}_{0}|^{4}$ in $W_{2}(\theta,\mathbf{II}(\mathbf{y}))$ scales this energy density so that it is invariant under conformal transformations of the unit cell, i.e., given by $(\tilde{\mathbf{u}}_{0},\tilde{\mathbf{v}}_{0})\rightarrow\lambda(\tilde{\mathbf{u}}_{0},\tilde{\mathbf{v}}_{0})$ and $(\mathbf{u}(\theta),\mathbf{v}(\theta))\rightarrow\lambda(\mathbf{u}(\theta),\mathbf{v}(\theta))$ for all $\lambda>0$. Finally, $L_{\Omega}>0$ denotes a characteristic lengthscale of the domain $\Omega$. It is introduced in $W_{2}$ so that $c_{2}$ has dimensions of energy per unit area, just like $c_{1}$.
@@
Note that, while we eventually anticipate fitting $c_{1},c_{2},d_{1},d_{2},d_{3}>0$ to experiments, our asymptotic analysis [64] suggests that these moduli should obey the relations
$c_{1}\gg c_{2}\gg d_{1}\sim d_{2}$ (23)
in the typical case. We also assumed $d_{3}\ll c_{1},\ldots,d_{2}$ in [64] to simplify the analysis. However, the appropriate choice of this folding modulus really depends in a delicate way on the stiffness ratio between the panels and folds. In Section 5, for instance, we choose $d_{3}$ to be larger than $d_{1,2}$ in most examples, as this choice produces simulations that conform well to the paper models of the origami in Fig. 1; $d_{3}$ should always be $\ll c_{1},c_{2}$.
As a final point on the constitutive model, we take the the ambient space on which to define and study this energy to be
@@ 3.2 Boundary conditions and applied forces
We investigate the effective behavior of parallelogram origami under Dirichlet boundary conditions, corresponding to applied deformation and slope, as well as applied forces, generalized moments, and distributed loads. Let $\Gamma_{\text{d}},\Gamma_{\text{n}}\subset\partial\Omega$ denote relatively open subsets of the boundary of $\Omega$. We prescribe the deformation and slope, respectively, via
@@
Let $\Gamma_{\text{t}}:=\partial\Omega\setminus\Gamma_{\text{d}}$ and $\Gamma_{\text{m}}:=\partial\Omega\setminus\Gamma_{\text{n}}$ denote the boundary sets on which we have not applied the deformation or slope. We allow for applied forces on these sets via the external energy potential
@@ 3.4 Weak formulation and equilibrium equations
This completes our description of the weak formulation of the equilibrium equations. In summary, if $(\mathbf{y},\theta)\in V_{\Gamma}$ is a local minimizer to the energy $E$, then (31) holds for all $(\mathbf{v},\eta)\in V_{0}$ for the generalized stresses $\mathbf{P},\mathcal{H},q,\mathbf{j}$ defined by (32-38).
@@ 4.1 Preliminaries
We always consider a discretization of the domain $\Omega$ via a shape-regular conforming triangulation $\mathcal{T}^{h}$ with $h_{T}:=\text{diam}(T)$ and $h:=\max_{T\in\mathcal{T}_{h}}h_{T}$ (see [9] for details). We denote by $\mathcal{E}^{h}$ the set of all edges of $\mathcal{T}^{h}$, $\mathcal{E}_{\Gamma}^{h}$ the set of all boundary edges $e\in\mathcal{E}^{h}$ such that $e\subset\partial\Omega$, and $\mathcal{E}_{\text{int}}^{h}:=\mathcal{E}^{h}\setminus\mathcal{E}_{\Gamma}^{h}$ the set of all internal edges. We also consider the subsets $\mathcal{E}_{\Gamma}^{h}$ denoted $\mathcal{E}_{\Gamma_{\text{d}}}^{h},\mathcal{E}_{\Gamma_{\text{t}}}^{h},\mathcal{E}_{\Gamma_{\text{n}}}^{h},\mathcal{E}_{\Gamma_{\text{m}}}^{h}$, which collect the edges on $\Gamma_{\text{d}}$, $\Gamma_{\text{t}},$ $\Gamma_{\text{n}}$ and $\Gamma_{\text{m}}$, respectively. For any interior edge $e\in\mathcal{E}_{\text{int}}^{h}$ shared by two triangles, denoted $T^{+}$ and $T^{-}$, we define $\mathbf{n}_{e}$ as the unit normal to $e$ that points from $T^{+}$ to $T^{-}$. As field variables defined on the triangulation $\mathcal{T}^{h}$ can jump across edges, we define the jump and average of a field variable ()($\cdot$)( ⋅ ) (any scalar, vector, or $n^{\text{th}}$-order tensor field) across the edge $e$ by $\llbracket(\cdot)\rrbracket_{e}:=(\cdot)^{+}-(\cdot)^{-}$ and {{()}}$\{\mskip-5.0mu\{(\cdot)\}\mskip-5.0mu\}_{e}:=\frac{1}{2}((\cdot)^{+}+(\cdot)^{-})$, respectively, where $(\cdot)^{\pm}$ denotes the limit of the restriction of the field variable $(\cdot)|_{T^{\pm}}$ to the edge $e$. For an exterior edge $e\in\mathcal{E}_{\Gamma}^{h}$, we take $\mathbf{n}_{e}\equiv\mathbf{n}$ to be the outward unit normal of $T$ on $e$. Finally, the length of any edge $e\in\mathcal{E}^{h}$ is given by $|e|$.
@@ 4.2 Weak formulation using the interior penalty method
$\mathscr{A}^{h}_{\text{con}}((\mathbf{y}_{h},\theta_{h});(\mathbf{w}_{h},\eta_{h})):=-\sum_{e\in\mathcal{E}_{\text{int}}^{h}}\int_{e}\big{(}\{\mskip-5.0mu\{\mathcal{H}(\mathbf{y}_{h},\theta_{h})\}\mskip-5.0mu\}_{e}\cdot\mathbf{n}_{e}\big{)}\colon\llbracket\nabla\mathbf{w}_{h}\rrbracket_{e}\mathop{}\!\mathrm{d}s-\int_{\Gamma_{\text{n}}}\big{(}\mathcal{H}(\mathbf{y}_{h},\theta_{h})\cdot\mathbf{n}\big{)}\colon\nabla\mathbf{w}_{h}\mathop{}\!\mathrm{d}s,$ (47)
@@
$\mathscr{A}^{h}_{\text{sta}}((\mathbf{y}_{h},\theta_{h});(\mathbf{w}_{h},\eta_{h})):=\sum_{e\in\mathcal{E}^{h}_{\text{int}}}\frac{\alpha}{|e|}\int_{e}\llbracket\nabla\mathbf{y}_{h}\rrbracket_{e}:\llbracket\nabla\mathbf{w}_{h}\rrbracket_{e}\mathop{}\!\mathrm{d}s.$ (48)
It corresponds to a jump energy of the form $\tfrac{1}{2}\sum_{e\in\mathcal{E}^{h}_{\text{int}}}\frac{\alpha}{|e|}\int_{e}\llbracket_{e}\nabla\mathbf{y}_{h}\rrbracket^{2}\mathop{}\!\mathrm{d}x$ and provides coercivity of the overall discrete energy for $\alpha>0$ large enough, ensuring the existence of a solution to (44). The final term $\mathscr{A}_{\text{bnd}}^{h}$ seeks to impose the slope boundary condition weakly via
@@ 5 Examples
This choice, while seemingly more cumbersome than, e.g., fixing a unit square domain, allows us to construct origami deformations of $M\times M$ cells that match the analytical and simulation results of the effective model.
@@ 5.1 Mechanism deformations
Perhaps the most interesting aspect of these plots is the stress-softening behavior. This type of behavior was first theoretically reported in [63] and subsequently reproduce in a bar and hinge model in [33]. We see from our curves that this loss of monotonicity occurs when $d_{3}/c_{1}$ is sufficiently small. The curves also quickly spring back up, a behavior that arises because the bulk energy $W_{1}(\theta,\mathbf{I}(\mathbf{y}))$ in (20-LABEL:eq:W12Def) is scaled by $1/\sqrt{\det\mathbf{I}(\mathbf{y})}$, which elastically penalizes deformations where the panels are close to overlapping one another. In particular, when the actuation approaches the fully folded states (see Fig. 6(vi) and (VI)), the second term in the stress $\mathbf{P}_{1}(\mathbf{y},\theta)$ in (LABEL:eq:P1P2) dominates the first, leading the force curves to veer upwards.
It is also interesting to compare and contrast the curves for Miura and Eggbox. We see that both sets of curves are qualitatively similar — the normlized force increases sub-linearly, before dipping a bit (depending on the ratio $d_{3}/c_{1})$ and then springing back up. However, most notably, the Miura pattern experiences significantly larger normalized force compared to the Eggbox one, even though the comparison is for the same underlying moduli parameters $c_{1}$ and $d_{3}$. This feature highlights the dominant role of geometry in this calculation. To get more granular, observe that the normalized stiffness of each pattern in its undeformed configuration is given by
@@ 5.2 Pure bending and twisting deformations
Although we eventually specialize to orthogonal parameterizations for $\mathbf{u}(\theta)$ and $\mathbf{v}(\theta)$ (see (50-51)) and rectangular domains (see (52)), we start fairly general here by assuming only that $\Omega$ is simply connected (has no holes). Then, using a variant on the Gauss and Codazzi-Mainardi compatibility conditions from differential geometry [17], the constraints in (60) can be recast equivalently as a coupled system of algebraic and differential equations: find a set of fields $(\bm{\omega}_{\mathbf{u}_{0}},\bm{\omega}_{\mathbf{v}_{0}},\theta)\colon\Omega\rightarrow\mathbb{R}^{3}\times\mathbb{R}^{3}\times(\theta^{-},\theta^{+})$ such that
$\begin{cases}\bm{\omega}_{\mathbf{u}_{0}}\times\mathbf{v}(\theta)+(\partial_{\mathbf{u}_{0}}\theta)\mathbf{v}^{\prime}(\theta)=\bm{\omega}_{\mathbf{v}_{0}}\times\mathbf{u}(\theta)+(\partial_{\mathbf{v}_{0}}\theta)\mathbf{u}^{\prime}(\theta)\\ \bm{\omega}_{\mathbf{u}_{0}}\cdot\mathbf{v}^{\prime}(\theta)=\bm{\omega}_{\mathbf{v}_{0}}\cdot\mathbf{u}^{\prime}(\theta)\\ \partial_{\mathbf{u}_{0}}\bm{\omega}_{\mathbf{v}_{0}}-\partial_{\mathbf{v}_{0}}\bm{\omega}_{\mathbf{u}_{0}}=\bm{\omega}_{\mathbf{v}_{0}}\times\bm{\omega}_{\mathbf{u}_{0}}.\end{cases}$ (61)
To explain the equivalence, there is first, for any solution to (61), a unique rotation field $\mathbf{R}\colon\Omega\rightarrow SO(3)$ that solves the differential equations $\partial_{\mathbf{u}_{0}}\mathbf{R}=\mathbf{R}(\bm{\omega}_{\mathbf{u}_{0}}\times)$ and $\partial_{\mathbf{v}_{0}}\mathbf{R}=\mathbf{R}(\bm{\omega}_{\mathbf{v}_{0}}\times)$ subject to $\mathbf{R}(\langle\mathbf{x}\rangle)=\mathbf{I}$, where $\langle\mathbf{x}\rangle=\frac{1}{|\Omega|}\int_{\Omega}\mathbf{x}\mathop{}\!\mathrm{d}x$ is the average point inside the domain $\Omega$. Furthermore, there is a unique deformation $\mathbf{y}\colon\Omega\rightarrow\mathbb{R}^{3}$ that solves the differential equations $\partial_{\mathbf{u}_{0}}\mathbf{y}=\mathbf{R}\mathbf{u}(\theta)$ and $\partial_{\mathbf{v}_{0}}\mathbf{y}=\mathbf{R}\mathbf{v}(\theta)$ subject to $\mathbf{y}(\langle\mathbf{x}\rangle)=\langle\mathbf{x}\rangle$. [^8] The specification of $\mathbf{R}$ and $\mathbf{y}$ at a point inside the domain, e.g., the average point $\langle\mathbf{x}\rangle$, guarantees uniqueness of the two fields. Without such specification, the fields are unique only up to a rigid body motion. The pair $(\mathbf{y},\theta)$, constructed as above, satisfies (60); moreover, the procedure to obtain $(\mathbf{y},\theta)$ based on a solution to (61) is exhaustive so long as we concern ourselves with sufficiently smooth solutions to (60).
Actually, the system in (61) can be further refined. Notice that the first two constraints are four linear equations in the six DOFs represented by the vectors $\bm{\omega}_{\mathbf{u}_{0}}$ and $\bm{\omega}_{\mathbf{v}_{0}}$. They are solved by writing
$\bm{\omega}_{\mathbf{u}_{0}}=\tau\mathbf{u}(\theta)+\kappa(\mathbf{u}^{\prime}(\theta)\cdot\mathbf{u}(\theta))\mathbf{v}(\theta)+\Big{(}\frac{(\partial_{\mathbf{u}_{0}}\theta)(\mathbf{v}^{\prime}(\theta)\cdot\mathbf{u}(\theta))-(\partial_{\mathbf{v}_{0}}\theta)(\mathbf{u}^{\prime}(\theta)\cdot\mathbf{u}(\theta))}{\mathbf{e}_{3}\cdot(\mathbf{u}(\theta)\times\mathbf{v}(\theta))}\Big{)}\mathbf{e}_{3},$ (62)
$\bm{\omega}_{\mathbf{v}_{0}}=\kappa(\mathbf{v}^{\prime}(\theta)\cdot\mathbf{v}(\theta))\mathbf{u}(\theta)-\tau\mathbf{v}(\theta)+\Big{(}\frac{(\partial_{\mathbf{u}_{0}}\theta)(\mathbf{v}^{\prime}(\theta)\cdot\mathbf{v}(\theta))-(\partial_{\mathbf{v}_{0}}\theta)(\mathbf{u}^{\prime}(\theta)\cdot\mathbf{v}(\theta))}{\mathbf{e}_{3}\cdot(\mathbf{u}(\theta)\times\mathbf{v}(\theta))}\Big{)}\mathbf{e}_{3},$
where $\kappa,\tau\colon\Omega\rightarrow\mathbb{R}$ are now the independent variables along with $\theta$. Substituting this parameterization into the final equation in (61) produces a system of three nonlinear PDEs in the three unknowns $\kappa,\tau$, and $\theta$ (a lengthy but explicit set of equations (140-142) in Section 5.1 of [64]).
@@
A parameterization of all possible solutions $(\kappa,\tau,\theta)$ to the PDE system seems out of reach. However, we can produce some illustrative examples based on a 1D ansatz. In the setting where the lattice vectors are orthogonal $\mathbf{u}(\theta)=\lambda_{1}(\theta)\mathbf{e}_{1}$ and $\mathbf{v}(\theta)=\lambda_{2}(\theta)\mathbf{e}_{2}$, like in the Miura and Eggbox patterns, we set
@@
where $\tilde{\mathbf{e}}_{1},\tilde{\mathbf{e}}_{2}$ are the standard basis vectors on $\mathbb{R}^{2}$. Substituting the ansatz in (63) into $(\kappa,\tau,\theta)$ PDE leads to a coupled set of ODEs in the scalar variable $s:=\mathbf{x}\cdot\lambda_{2}^{-1}(0)\tilde{\mathbf{e}}_{2}$ for the fields $(\theta_{\text{1D}}(s),\kappa_{\text{1D}}(s),\tau_{\text{1D}}(s))$, which we solved in Appendix C [64]. The general solutions are
$\kappa_{\text{1D}}=\frac{c_{\kappa}}{\lambda_{1}(\theta_{\text{1D}})\lambda_{1}^{\prime}(\theta_{\text{1D}})\lambda_{2}^{2}(\theta_{\text{1D}})},\quad\tau_{\text{1D}}=\frac{c_{\tau}}{\lambda^{2}_{1}(\theta_{\text{1D}})}$ (64)
@@
subject to the initial conditions $\theta_{\text{1D}}(0)=\overline{\theta}$ and $\theta_{\text{1D}}^{\prime}(0)=\overline{\zeta}$.
In this parameterization, $\kappa_{\text{1D}}$ represents bending while $\tau_{\text{1D}}$ represents twist. As always, $\theta_{\text{1D}}$ represents cell-by-cell actuation. Thus, pure bending modes are obtained by solving (64) and (65) with $c_{\tau}=0$; pure twisting modes are obtained likewise by solving these equations with $c_{\kappa}=0$.
Fig. 7 and 8 compare analytical solutions of pure bending and twisting modes for both Miura and Eggbox to analogous simulations using our model and numerical method. In all cases, the $\theta$-dependence is specified by (50) for Miura and (51) for Eggbox and $\Omega$ is as in (52). Concerning the analytics on Miura, the inputs to the ODE in (65) are $(c_{\kappa},c_{\tau},\theta_{\text{1D}}(0),\theta_{\text{1D}}^{\prime}(0))=(1.518,0,-\pi/6+0.2,16)$ for pure bending and $=(0,1.5205,-\pi/6+0.2,11)$ for pure twisting; the analogous inputs for Eggbox are $=(1.43,0,0.4,-1.5)$ and $=(0,1.9,-0.5,4)$ for bend and twist, respectively. These inputs are chosen so that the solution $\theta_{\text{1D}}(s)$ is approximately symmetric about $s=1/2$, corresponding to the midline of the domain $\Omega$ in the $\mathbf{v}_{0}$ direction. After obtaining the solutions, we substitute $(\theta_{\text{1D}},\kappa_{\text{1D}},\tau_{\text{1D}})\equiv(\theta,\kappa,\tau)$ into (LABEL:eq:theOmegaParams). The rotation field $\mathbf{R}$ and deformation $\mathbf{y}$ are then determined by solving the differential equations $\partial_{\mathbf{u}_{0}}\mathbf{R}=\mathbf{R}(\bm{\omega}_{\mathbf{u}_{0}}\times)$, $\partial_{\mathbf{v}_{0}}\mathbf{R}=\mathbf{R}(\bm{\omega}_{\mathbf{v}_{0}}\times)$, $\partial_{\mathbf{u}_{0}}\mathbf{y}=\mathbf{R}\mathbf{u}(\theta),$ and $\partial_{\mathbf{v}_{0}}\mathbf{y}=\mathbf{R}\mathbf{v}(\theta)$, as explained in Appendix C of [64]. Finally, Figs. 7, 8(a) and (e) show plots of origami deformations constructed from these solutions using an ansatz outlined in Appendix F. Notice that the Miura takes the shape of a saddle for pure bending $(c_{\tau}=0)$, while Eggbox takes that of the opposite bending profile in a cap. In contrast, the pure twist deformations of these pattern $(c_{\kappa}=0)$ are qualitatively similar.

### v3

@@ 3 Modeling framework
Building on the theory, we now develop a constitutive model well-adapted to simulate the effective behavior of soft modes in parallelogram origami. A basic challenge is that it is hard to numerically implement exact geometrical constraints, like the ones in $A_{\text{eff}}$. This is especially true under prototypical displacement and traction boundary conditions. We seek instead a practical continuum constitutive model that is faithful to the theory but also efficient to simulate and predictive.
@@ 3.1 Constitutive model
The division by $\sqrt{\det\mathbf{I}(\mathbf{y})}$ in the $W_{1}(\theta,\mathbf{I}(\mathbf{y}))$ seeks to prevent interpenetration of matter in the model, while the division by $|\tilde{\mathbf{u}}_{0}|^{4}|\tilde{\mathbf{v}}_{0}|^{4}$ in $W_{2}(\theta,\mathbf{II}(\mathbf{y}))$ scales this energy density so that it is invariant under conformal transformations of the unit cell given by $(\tilde{\mathbf{u}}_{0},\tilde{\mathbf{v}}_{0})\rightarrow\lambda(\tilde{\mathbf{u}}_{0},\tilde{\mathbf{v}}_{0})$ and $(\mathbf{u}(\theta),\mathbf{v}(\theta))\rightarrow\lambda(\mathbf{u}(\theta),\mathbf{v}(\theta))$ for all $\lambda>0$. Finally, $L_{\Omega}>0$ denotes a characteristic lengthscale of the domain $\Omega$. It is introduced in $W_{2}$ so that $c_{2}$ has dimensions of energy per unit area, just like $c_{1}$.
@@
[…]
$c_{1}\sim\mu,\quad c_{2}\sim\mu\ell^{2},\quad d_{1}\sim d_{2}\sim\mu\Big{(}\frac{h}{L_{\Omega}}\Big{)}^{2},\quad d_{3}\sim\mu_{\text{f}}\frac{w}{L_{\Omega}}\ell^{-1}.$ (25)
In practice, we anticipate fitting $c_{1},c_{2},d_{1},d_{2},d_{3}$ to experiments, rather than being too beholden to these scaling laws. Nevertheless, some general features are worth noting. Since an origami metamaterial satisfies $h/L_{\Omega}\ll\ell\ll 1$ in the typical case, these scaling laws suggest that these moduli should obey the relations
$c_{1}\gg c_{2}\gg d_{1}\sim d_{2}.$ (26)
Notably absent in these relations is $d_{3}$. As a practical matter, we studied the regime of very soft folds in [64] (corresponding to $d_{3}\ll c_{1},\ldots,d_{2}$) to derive the plate theory in (19) by asymptotic analysis. However, the physically appropriate choice for this modulus depends delicately on the properties of the panels and folds. In general, we expect $\mu_{\text{f}}\ll\mu$ because the folds are often made by perforating and/or plastically yielding the creases or are simply made of a more compliant material than the panels. In addition, $w$ typically satisfies $w/L_{\Omega}\sim h/L_{\Omega}\ll\ell$. So we can conclude that $d_{3}\ll c_{1}$; however, its size relative to $c_{2},d_{1},d_{2}$ is not universal. In Section 5, for instance, we choose $d_{3}$ to be larger than $d_{1,2}$ and smaller than $c_{2}$ in most examples, as this choice produces simulations that conform well to the paper models of the origami in Fig. 1.
As a final point on the constitutive model, we take the ambient space on which to define and study this energy to be
@@ 3.2 Boundary conditions and applied forces
We investigate the effective behavior of parallelogram origami under Dirichlet boundary conditions, corresponding to prescribed deformation and slope, as well as applied forces, generalized moments, and distributed loads. Let $\Gamma_{\text{d}},\Gamma_{\text{n}}\subset\partial\Omega$ denote relatively open subsets of the boundary of $\Omega$. We prescribe the deformation and slope, respectively, via
@@
Let $\Gamma_{\text{t}}:=\partial\Omega\setminus\Gamma_{\text{d}}$ and $\Gamma_{\text{m}}:=\partial\Omega\setminus\Gamma_{\text{n}}$ denote the boundary sets on which we have not prescribed the deformation or slope. We allow for applied forces on these sets via the external energy potential
@@ 3.4 Weak formulation and equilibrium equations
This completes our description of the weak formulation of the equilibrium equations. In summary, if $(\mathbf{y},\theta)\in V_{\Gamma}$ is a local minimizer to the energy $E$, then (34) holds for all $(\mathbf{v},\eta)\in V_{0}$ for the generalized stresses $\mathbf{P},\mathcal{H},q,\mathbf{j}$ defined by (35-41).
@@ 4.1 Preliminaries
We always consider a discretization of the domain $\Omega$ via a shape-regular conforming triangulation $\mathcal{T}^{h}$ with $h_{T}:=\text{diam}(T)$ and $h:=\max_{T\in\mathcal{T}_{h}}h_{T}$ (see [9] for details). We denote by $\mathcal{E}^{h}$ the set of all edges of $\mathcal{T}^{h}$, $\mathcal{E}_{\Gamma}^{h}$ the set of all boundary edges $e\in\mathcal{E}^{h}$ such that $e\subset\partial\Omega$, and $\mathcal{E}_{\text{int}}^{h}:=\mathcal{E}^{h}\setminus\mathcal{E}_{\Gamma}^{h}$ the set of all internal edges. We also consider the subsets $\mathcal{E}_{\Gamma}^{h}$ denoted $\mathcal{E}_{\Gamma_{\text{d}}}^{h},\mathcal{E}_{\Gamma_{\text{t}}}^{h},\mathcal{E}_{\Gamma_{\text{n}}}^{h},\mathcal{E}_{\Gamma_{\text{m}}}^{h}$, which collect the edges on $\Gamma_{\text{d}}$, $\Gamma_{\text{t}},$ $\Gamma_{\text{n}}$ and $\Gamma_{\text{m}}$, respectively. For any interior edge $e\in\mathcal{E}_{\text{int}}^{h}$ shared by two triangles, denoted $T^{+}$ and $T^{-}$, we define $\mathbf{n}_{e}$ as the unit normal to $e$ that points from $T^{+}$ to $T^{-}$. As field variables defined on the triangulation $\mathcal{T}^{h}$ can jump across edges, we define the jump and average of a field variable ()($\cdot$)( ⋅ ) (any scalar, vector, or $n^{\text{th}}$-order tensor field) across the edge $e$ by $\llbracket(\cdot)\rrbracket:=(\cdot)^{+}-(\cdot)^{-}$ and {{()}}$\{\mskip-5.0mu\{(\cdot)\}\mskip-5.0mu\}:=\frac{1}{2}((\cdot)^{+}+(\cdot)^{-})$, respectively, where $(\cdot)^{\pm}$ denotes the limit of the restriction of the field variable $(\cdot)|_{T^{\pm}}$ to the edge $e$. For an exterior edge $e\in\mathcal{E}_{\Gamma}^{h}$, we take $\mathbf{n}_{e}\equiv\mathbf{n}$ to be the outward unit normal of $T$ on $e$. Finally, the length of any edge $e\in\mathcal{E}^{h}$ is given by $|e|$.
@@ 4.2 Weak formulation using the interior penalty method
$\mathscr{A}^{h}_{\text{con}}((\mathbf{y}_{h},\theta_{h});(\mathbf{w}_{h},\eta_{h})):=-\sum_{e\in\mathcal{E}_{\text{int}}^{h}}\int_{e}\big{(}\{\mskip-5.0mu\{\mathcal{H}(\mathbf{y}_{h},\theta_{h})\}\mskip-5.0mu\}\cdot\mathbf{n}_{e}\big{)}\colon\llbracket\nabla\mathbf{w}_{h}\rrbracket\mathop{}\!\mathrm{d}s-\int_{\Gamma_{\text{n}}}\big{(}\mathcal{H}(\mathbf{y}_{h},\theta_{h})\cdot\mathbf{n}\big{)}\colon\nabla\mathbf{w}_{h}\mathop{}\!\mathrm{d}s,$ (50)
@@
$\mathscr{A}^{h}_{\text{sta}}((\mathbf{y}_{h},\theta_{h});(\mathbf{w}_{h},\eta_{h})):=\sum_{e\in\mathcal{E}^{h}_{\text{int}}}\frac{\alpha}{|e|}\int_{e}\llbracket\nabla\mathbf{y}_{h}\rrbracket:\llbracket\nabla\mathbf{w}_{h}\rrbracket\mathop{}\!\mathrm{d}s.$ (51)
It corresponds to a jump energy of the form $\tfrac{1}{2}\sum_{e\in\mathcal{E}^{h}_{\text{int}}}\frac{\alpha}{|e|}\int_{e}\llbracket\nabla\mathbf{y}_{h}\rrbracket^{2}\mathop{}\!\mathrm{d}s$ and provides coercivity of the overall discrete energy for $\alpha>0$ large enough, ensuring the existence of a solution to (47). The final term $\mathscr{A}_{\text{bnd}}^{h}$ seeks to impose the slope boundary condition weakly via
@@ 5 Examples
This choice, while seemingly more cumbersome than fixing a unit square domain, ensures that the aspect ratio of $\Omega$ is the same as that of a unit cell of the origami. As such, we can compare analytical and simulation results of the effective model directly to origami deformations with a discrete $M\times M$ number of cells. We do such a comparison for all examples in sequel. The origami deformations are constructed to match that of the effective model by the method outlined in Appendix F.
@@ 5.1 Mechanism deformations
Perhaps the most interesting aspect of these plots is the stress-softening behavior. This type of behavior was first theoretically reported in [63] and subsequently reproduced in a bar and hinge model in [33]. We see from our curves that this loss of monotonicity occurs when $d_{3}/c_{1}$ is sufficiently small. The curves also quickly spring back up, a behavior that arises because the bulk energy $W_{1}(\theta,\mathbf{I}(\mathbf{y}))$ in (22-LABEL:eq:W12Def) is scaled by $1/\sqrt{\det\mathbf{I}(\mathbf{y})}$, which elastically penalizes deformations where the panels are close to overlapping one another. In particular, when the actuation approaches the fully folded states (see Fig. 6(vi) and (VI)), the second term in the stress $\mathbf{P}_{1}(\mathbf{y},\theta)$ in (LABEL:eq:P1P2) dominates the first, leading the force curves to veer upwards.
It is also interesting to compare and contrast the curves for Miura and Eggbox. We see that both sets of curves are qualitatively similar — the normalized force increases sub-linearly, before dipping a bit (depending on the ratio $d_{3}/c_{1})$ and then springing back up. However, most notably, the Miura pattern experiences significantly larger normalized force compared to the Eggbox one, even though the comparison is for the same underlying moduli parameters $c_{1}$ and $d_{3}$. This feature highlights the dominant role of geometry in this calculation. To get more granular, observe that the normalized stiffness of each pattern in its undeformed configuration is given by
@@ 5.2 Pure bending and twisting deformations
Let’s focus on the setting of orthogonal Bravais lattice vectors $\mathbf{u}(\theta)=\lambda_{1}(\theta)\mathbf{e}_{1}$ and $\mathbf{v}(\theta)=\lambda_{2}(\theta)\mathbf{e}_{2}$, consistent with the Miura and Eggbox examples. In such cases, the constraints in (63) are equivalent to parameterizing the first and second fundamental forms as
$\mathbf{I}(\mathbf{y})=\begin{pmatrix}\tfrac{\lambda^{2}_{1}(\theta)}{\lambda_{1}^{2}(0)}&0\\ 0&\tfrac{\lambda^{2}_{2}(\theta)}{\lambda_{2}^{2}(0)}\end{pmatrix},\quad\mathbf{II}(\mathbf{y})=\begin{pmatrix}-\tfrac{\lambda_{1}^{\prime}(\theta)}{\lambda_{1}(\theta)}\kappa&\tau\\ \tau&\tfrac{\lambda_{2}^{\prime}(\theta)}{\lambda_{2}(\theta)}\kappa\end{pmatrix}$ (64)
for compatible curvature fields $\kappa,\tau\colon\Omega\rightarrow\mathbb{R}$. Here $\kappa$ describes a bending curvature, while $\tau$ represents twist. As always, $\theta\colon\Omega\rightarrow(\theta^{-},\theta^{+})$ is the cell-by-cell actuation. Of course, these fields cannot be arbitrary. Much like the statement that a deformation gradient must be curl-free to admit a deformation, the metric and curvature tensors in (64) must solve the Gauss and Codazzi-Mainardi equations from differential geometry [17]. This produces a system of three nonlinear PDEs in the three unknowns $\kappa,\tau$, and $\theta$, fully characterizing the existence of a deformation $\mathbf{y}$ solving (64).
We do not write out the full PDE system here because it is cumbersome and we do not know how to solve it in general. Instead we produce some illustrative examples based on a 1D ansatz. By setting
@@
the PDEs becomes a coupled set of ODEs in the scalar variable $s:=\mathbf{x}\cdot\lambda_{2}^{-1}(0)\tilde{\mathbf{e}}_{2}$ for the fields $(\theta_{\text{1D}}(s),\kappa_{\text{1D}}(s),\tau_{\text{1D}}(s))$, which we solved in Appendix C [64]. The general solutions are
$\kappa_{\text{1D}}=\frac{c_{\kappa}}{\lambda_{1}^{\prime}(\theta_{\text{1D}})\lambda_{2}(\theta_{\text{1D}})},\quad\tau_{\text{1D}}=\frac{c_{\tau}}{\lambda^{2}_{1}(\theta_{\text{1D}})}$ (66)
@@
subject to the initial conditions $\theta_{\text{1D}}(0)=\overline{\theta}$ and $\theta_{\text{1D}}^{\prime}(0)=\overline{\zeta}$. Note that pure bending modes in the parameterization are obtained by solving (66) and (67) with $c_{\tau}=0$; pure twisting modes are obtained likewise by solving these equations with $c_{\kappa}=0$.
Fig. 7 and 8 compare analytical solutions of pure bending and twisting modes for both Miura and Eggbox to analogous simulations using our model and numerical method. In all cases, the $\theta$-dependence is specified by (53) for Miura and (54) for Eggbox and $\Omega$ is as in (55). Concerning the analytics on Miura, the inputs to the ODE in (67) are $(c_{\kappa},c_{\tau},\theta_{\text{1D}}(0),\theta_{\text{1D}}^{\prime}(0))=(1.518,0,-\pi/6+0.2,16)$ for pure bending and $=(0,1.5205,-\pi/6+0.2,11)$ for pure twisting; the analogous inputs for Eggbox are $=(1.43,0,0.4,-1.5)$ and $=(0,1.9,-0.5,4)$ for bend and twist, respectively. These inputs are chosen so that the solution $\theta_{\text{1D}}(s)$ is approximately symmetric about $s=1/2$, corresponding to the midline of the domain $\Omega$ in the $\mathbf{v}_{0}$ direction. After obtaining the solutions, we substitute $(\theta_{\text{1D}},\kappa_{\text{1D}},\tau_{\text{1D}})\equiv(\theta,\kappa,\tau)$ into (64). The deformation $\mathbf{y}$ is then constructed from these compatible fields using a procedure outlined at the end of Appendix C of [64]. Finally, Figs. 7, 8(a) and (e) show plots of the corresponding origami deformations built from these solutions using the ansatz in Appendix F. Notice that the Miura takes the shape of a saddle for pure bending $(c_{\tau}=0)$, while Eggbox takes that of the opposite bending profile in a cap. In contrast, the pure twist deformations of these pattern $(c_{\kappa}=0)$ are qualitatively similar.